A light brew coffee sells for $12 per pound is to be mixed with a dark brew coffee that sells for $9 per pound to get a 5-pound mixture that will sell for $10 per pound. How many pounds of light brew and dark brew coffee should be used?
Answer by macston(5194)

.
T=amount of $12 coffee; N=amount of $9 coffee.
N+T=5lbs
N=5lbs-T Use this to substitute for N
$9N+$12T=$10(5lbs) Substitute for N from above.
$9(5lbs-T)+$12T=$50lbs
$45lbs-$9T+$12T=$50lbs Subtract $45lbs from each side.
$3T=$5lbs Divide each side by $3.
T=5/3 lbs ANSWER 1:5/3 lbs of light brew ($12) coffee should be used.
N=5lbs-T=5lbs-5/3lbs=10/3 lbs ANSWER 10/3 lbs of dark brew ($9) should be used.
.
CHECK:
.
$9N+$12T=$10(5lbs)
$9(10/3)+$12(5/3)=$10(5lbs)
$90/3+$60/3=$50
$30+$20=$50
$50=$50